A quarterback throws a football to a stationary receiver 31.5m away from him. If the football is thrown at an initial angle of 40° to the ground, at what initial speed must the quarterback throw the ball for it to reach the receiver?

1
Expert's answer
2022-05-30T15:18:45-0400
"R=\frac{v_0^2\sin2\theta}{g}"

"v_0=\sqrt{gR/\sin2\theta}\\
=\sqrt{9.8*31.5/\sin80^\circ}=313\:\rm m/s"